OpenStudy (anonymous):

Help Please!! Find the equation of the line in slope-intercept form that is perpendicular to y-2=1/4(x-8) and passes through (-12,8)

OpenStudy (abmon98):

If two lines are perpendicular, their slopes are negative reciprocals. m1*m2=-1 y-2=1/4(x-8) y-y1=m(x-x1) m1=1/4 1/4*m2=-1 m2=-1*4/1=-4 y-8=-4(x--12)

OpenStudy (anonymous):

thank you @Abmon98

OpenStudy (abmon98):

your welcome